# Beyond Single-Turn: A Survey on Multi-Turn Interactions with Large Language Models

### v4

@@ 4.2.2 Topic Switching & Discontinuous Reasoning
Real conversations often shift topics or return to earlier subjects after a detour. Such discontinuities pose difficulties for LLMs, which may either wrongly carry over context from the previous topic or fail to recall the relevant earlier context when a topic is resumed. For example, after discussing topic A, then briefly B, a question about A again might confuse the model or lead to a non-sequitur response. Current models have trouble with this kind of context management – they do not truly pause and resume threads of conversation like humans. Instead, models either treat each turn in isolation (losing the prior thread) or conflate unrelated context. Effective attention re-allocation is required for handling topic switches, yet remains unsolved. Recent benchmarks like MT-Bench [21] and MT-Bench-101 [30] contain category shifts to test this, and results indicate that even top-tier models often break coherence when the conversation flow is non-linear. Better mechanisms for tracking multiple conversation topics and context segmentation are needed.
@@ 4.2.3 Proactive Information Seeking
In complex interactive settings such as medical diagnosis, troubleshooting, or tutoring, an ideal conversational agent should ask follow-up questions and guide the dialogue to gather missing information. Today’s general-purpose LLMs, however, are largely reactive — they answer questions when asked but rarely volunteer clarifying or exploratory queries unless explicitly prompted. This limits their effectiveness in diagnostic and problem-solving dialogues, where the burden of inquiry should not fall solely on the user. One challenge is that LLMs have been primarily trained on single-turn Q&A or straightforward instruction-following, so they lack exposure to multi-turn dialogue policies (the strategy of when to ask, when to inform) [192]. Data for such scenarios is scarce, and models tend to default to giving an immediate answer even if the query is underspecified or potentially incomplete. For instance, in a medical context, a user might say “I feel sick,” and a good agent would ask about specific symptoms; many LLMs instead try to provide a diagnosis or generic advice without sufficient clarification. Some recent work attempts to teach LLMs to clarify and confirm (e.g. asking follow-ups in ambiguous text-to-SQL tasks [192], but robust proactive dialogue behavior in arbitrary settings (medical, legal, technical diagnostics) is still an open challenge.
@@ 4.2.4 Multilingual & Code-Switching Scenarios
Multi-turn interactions that involve multiple languages (or mixing languages) introduce additional complexity. An LLM may handle a single-turn query in a non-English language reasonably well, yet maintaining context across languages in a conversation is far more difficult. Code-switching – where users alternate between languages in successive turns or even within a turn – often confuses models, leading to incorrect translations or lost context. For example, a user might ask a question in English, then follow up in Spanish; current models might not consistently link the follow-up to the previous question or could respond in the wrong language. Moreover, evaluation of multi-turn dialogue abilities has been heavily skewed toward English, and many models show performance drop-offs in less dominant languages. Cultural and linguistic knowledge may not transfer seamlessly across turns. There are also safety implications: mixing languages can bypass certain content filters or exploit weaknesses in a model’s multilingual alignment. Yoo et al. (2024) [263] found that code-switching in prompts can elicit undesirable behaviors from LLMs that would not surface with English-only inputs, revealing gaps in multilingual understanding and alignment. This underscores that handling multi-turn dialogues in a culturally and linguistically robust way is still an open challenge.
@@ 4.3.1 Dynamic Preference & Objective Adaptation
Unlike human assistants, LLMs have no true long-term learning or personalization within a conversation – they rely only on the provided context window. Adapting to a user’s preferences, tone, or goals over the course of an interaction is thus difficult. Current chat models will follow explicit style instructions (e.g. “please reply formally”) but they often miss subtler cues or changes in user preferences over time. For example, if a user seems confused, an ideal assistant might proactively simplify its language in subsequent turns; today’s LLM might continue at the same level unless directly told to adjust. Personalization is largely constrained to what the user explicitly repeats every turn. Developing dynamic adaptation is challenging because it requires the model to infer and remember user preferences or context that persist across turns (or even across sessions) without explicit reminders. Some initial research has explored allowing models to update a pseudo-persona or profile as they converse, but ensuring this happens reliably and safely is unresolved. The key difficulty is that the model’s parameters are fixed during inference – any adaptation must come from processing the conversation context itself. This leaves LLMs prone to either forgetting user preferences or overshooting (applying a style inappropriately broadly). Effective techniques for on-the-fly adaptation (short of fine-tuning) are still in early stages.
@@ 4.3.2 Knowledge Adaptation
Human conversation partners learn and adapt during dialogue – for instance, absorbing new facts the user provides or adjusting to corrections. Current LLMs, in contrast, have a fixed knowledge base at deployment and lack the ability to truly update their beliefs or knowledge states during a conversation. If a user informs the chatbot of a new piece of information in turn 5, the model might use it temporarily in subsequent responses (since it resides in the dialogue context), but this information isn’t incorporated into any lasting memory. Once the session resets, the “learning” is lost. Enabling continual learning or persistent memory in interactive settings is an important open challenge [264]. Approaches like retrieval-augmented generation partially address this by allowing the model to fetch relevant facts from an external database, but they still rely on pre-established knowledge sources. True on-the-fly learning would mean the model can update its internal representations based on user-provided data or feedback, without retraining from scratch. This is difficult due to the risk of catastrophic forgetting (updating a neural model on new data can degrade prior knowledge) and the danger of model misuse (users could insert false or malicious information). Some recent studies have explored using external memory modules or dynamic context augmentation as a surrogate for learning, effectively writing important new facts to a scratchpad that persists through the dialogue. Nonetheless, achieving a good balance between plasticity (learning new information) and stability (not overriding established correct knowledge or safety protocols) in real-time is an unsolved problem.
@@ 4.3.3 Robustness to Misinformation & Adversarial Inputs
In multi-turn settings, users may intentionally or unintentionally introduce misleading information, adversarial prompts, or attempts to derail the model (so-called jailbreaks). A major challenge is making LLMs resilient to these tactics. Over multiple turns, a malicious user can gradually manipulate the context or employ social-engineering style approaches to trick the model into breaking rules. Recent studies demonstrate that even top-tier aligned models (including GPT-4) can be gradually coerced into unsafe or policy-violating outputs through clever multi-turn strategies [265]. Attackers obscure harmful intents across several turns and lead the model astray with fabricated context, succeeding where a single-turn attack would fail. In other cases, a user might feed the model subtly incorrect facts each turn; the model typically lacks the fact-checking capability to catch these and will propagate the falsehoods. There is also the issue of prompt leakage – models revealing hidden system instructions or private content when pressed across turns (as discussed in the context of privacy below). All these adversarial scenarios highlight that sustained interactions open new vectors for exploitation. Current LLMs do not have robust defenses against multi-turn manipulation, beyond static safety training which can be sidestepped when the attack is staged gradually. Building conversational agents that can detect inconsistent or malicious user input and respond safely (or refuse) without derailing the interaction remains an open research problem.
@@ 4.4.1 Scalable Data Curation
Achieving scalable data collection for multi-turn interactions remains a significant challenge at the intersection of data engineering and model training. The community continues to explore a range of promising approaches, including leveraging domain-specific real dialogues, establishing collaborations to access new data sources, and developing innovative pipelines for synthetic data generation. Despite these efforts, scalability in curating conversational datasets is still an open and pressing issue.
@@
As we discussed in §2.1.1 and §2.2.2, as dataset sizes continue to grow, ensuring scalable, high-quality conversational data has proven crucial yet remains challenging. Among the domains we’ve examined, healthcare uniquely benefits from access to extensive real-world multi-turn, patient-doctor conversational datasets. These datasets, derived directly from authentic patient-doctor interactions, inherently provide rich conversational diversity and realism, making them highly suitable for continuous pre-training, SFT, and RLHF. However, replicating such high-quality data collection in other domains is challenging. To meet the rising demand for specialized conversational data across various fields, increased collaboration is required among researchers, governmental bodies, and institutional organizations to develop scalable systems for data collection. Unlike healthcare, most other domains lack established protocols for recording and anonymizing real-world conversations, representing a significant obstacle in dataset development.
Addressing these limitations, recent research has explored strategies for enriching and expanding existing datasets. For example, single-turn or limited-turn interactions can be converted into rich multi-turn dialogues either manually or via automated frameworks. Notable advancements include specialized methods and frameworks [120, 266, 165], as well as synthetic generation and rewriting techniques leveraging LLMs [122, 81, 32, 267]. However, these methods still fall short of ideal performance, and multiple studies have documented the drawbacks of synthetic data [268, 269, 270]. Even the most sophisticated scalable data approaches suffer from inherent limitations: they often lack the unexpected nuances and diversity of real-world conversations, risk compounding errors (model-generated dialogues might amplify factual inaccuracies or unnatural patterns if not monitored), introduce subtle inconsistencies, fail to capture genuine user spontaneity, or overrely on generic conversational patterns.
Going forward, research must focus on not just increasing the volume of multi-turn data, but doing so in a way that maintains or improves quality. This includes developing better automated dialogue generation that mimics human conversational nuances, creating robust filtering and refinement processes to catch errors in synthetic data, and establishing shared repositories of multi-turn dialogues that cover a broad range of domains and interaction styles. Only through such comprehensive efforts can we hope to meet the data demands of ever more sophisticated multi-turn LLM systems.
4.4.2 Metric Designing
@@
Beyond datasets, the metrics and criteria for judging multi-turn interactions remain open for improvement. Traditional metrics (e.g. single-turn accuracy or BLEU for responses) fail to capture the nuanced qualities of a dialogue. There is a need for more fine-grained metrics that evaluate specific aspects of multi-turn performance, as well as holistic metrics for entire conversations:
Capability Evaluation at A Fine-Grained Level
@@
Multi-turn benchmarks like MT-Bench-101 emphasize evaluating discrete abilities (e.g. logical consistency, factual recall, politeness) at the turn level [30]. Fine-grained scoring can reveal which skills degrade over a conversation (for example, a model might maintain grammar and politeness but lose factual accuracy after many turns). However, designing such granular rubrics is challenging – it requires identifying and weighting many sub-skills, and often necessitates expert annotation or LLM-as-judge for each aspect. The field is exploring rubrics and checklists for each turn or response (as in MultiChallenge’s instance-level rubrics [72], but no standardized set of fine-grained metrics has been universally adopted yet.
Evaluation on Long-Term Effectiveness
@@
Evaluating the outcome or global quality of a multi-turn exchange is still an open problem. Ideally, metrics should reflect whether the conversation as a whole was successful (e.g. the user’s goal achieved, or the model remained helpful and coherent throughout). This is hard to reduce to a single number. Some works look at conversational return-on-investment, measuring if additional turns are actually helping or just causing more confusion [27]. Others consider memory retention tests across turns or consistency checks at different conversation lengths. Still, we lack metrics for “did the model sustain high performance across 10+ turns?” or “did the conversation eventually converge to a correct/ useful outcome?”. Developing evaluation measures for long-range, interactive effectiveness (perhaps analogous to task success rates in dialogue systems) is an ongoing research area.
Cultural & Sociolinguistic Diversity
@@
Most evaluation setups to date focus on a narrow band of interaction styles (often Western, English-centric dialogues). This limits our ability to assess model performance for diverse users. A challenge is to create evaluation metrics and scenarios that account for cultural differences in conversation, multilingual nuances (as noted earlier), and dialect or style variations. For example, a model might perform well on a formal Q&A dialogue, but fare poorly on a casual, code-mixed conversation or when the user employs idiomatic expressions from a particular culture. Incorporating such diversity into benchmarks is crucial for a fair assessment. FairMT-Bench is a step in this direction, examining bias and fairness across demographic attributes in multi-turn settings [37]. The open problem is defining metrics that can quantify model behavior across a wide sociolinguistic spectrum – possibly by measuring bias, respectfulness, or user satisfaction for different user profiles. Ensuring our evaluation metrics are culturally inclusive and robust to different communication styles remains an important challenge.
@@
The evaluation of multi-turn LLM dialogues has increasingly turned to LLM-as-a-judge frameworks, where a strong model (e.g. GPT-4) assesses response quality in place of human annotators. Recent benchmarks like MT-Bench and platforms such as Chatbot Arena employ LLM judges to rank chatbot responses, achieving evaluation at scale with high consistency and low cost [21]. Indeed, studies have found that GPT-4 based evaluators can align with human preferences roughly 80% of the time, approaching inter-annotator agreement levels [21]. However, replacing human judgment introduces notable biases and limitations: an AI evaluator may display self-enhancement bias (preferring answers it or a similar model generated) and verbosity bias (rewarding unnecessarily long responses), among other systematic errors [21]. Automated judges can also struggle with factual correctness and nuanced context - for instance, they might overlook subtle flaws or value judgments that a human reviewer would catch [271]. To mitigate these issues, researchers have proposed hybrid approaches that combine LLM-based scoring with human oversight. Examples include prompting AI judges with explicit rubrics or checklists to guide their evaluations, and incorporating periodic human-in-the-loop audits of model ratings. This way, the field hopes to leverage the scalability of LLM evaluators while maintaining reliability and fairness close to human standards.
@@ 4.5.1 Bias Amplification
Multi-turn dialogues can inadvertently magnify a model’s biases with each turn. If a biased assumption slips into an early response, subsequent interactions might build on it, reinforcing stereotypes or unfair perceptions. Furthermore, the interactive nature of dialogue can lead the model to adapt to the user’s biases, resulting in amplified problematic content. Recent work on fairness in conversational AI underscores this risk. Fan et al. (2024) [37] find that LLMs are more prone to generating biased or discriminatory responses in multi-turn settings, showing greater bias accumulation than in single-turn prompts. For example, a slight gender bias in a first answer can become more extreme if the conversation continues in that vein. FairMT-Bench specifically evaluates such scenarios and reveals significant variability in how different models handle them – many state-of-the-art models show degraded fairness when faced with back-and-forth discussions involving sensitive attributes [37]. This challenge calls for improved bias mitigation techniques that operate across the dialogue timeline. It may not be enough to detoxify single responses; models might need mechanisms to self-monitor and adjust if they find their outputs drifting toward bias during a conversation. Additionally, data augmentation with diverse user profiles and enforcing consistency with ethical guidelines over multiple turns are potential avenues to address this. Bias in multi-turn AI interactions is an ethical concern with direct user impact, and it remains open how best to evaluate and reduce it without stifling free-flowing conversation.
@@ 4.5.2 Privacy Leakage
Prolonged conversations increase the risk of LLMs revealing sensitive information, either about the user or memorized from training data. One concern is that as a dialogue progresses, a user might share personal details that the model could later inadvertently expose or use inappropriately. Another well-documented issue is model memorization: large models sometimes recall specific training examples (such as private facts or copyrighted text) and may output them given the right prompts. Nasr et al. (2023) [272] demonstrated that LLMs can memorize chunks of their training data (like personal phone numbers, addresses, or secrets) and that adversaries can craft sequences of prompts to extract this hidden information. Multi-turn interactions give an attacker more opportunities to perform such extraction gradually, by establishing context and probing iteratively. Even if a single-turn query doesn’t trigger a leak, a series of cleverly steered queries might corner the model into revealing something it “knows”. Another aspect of privacy is system prompt or context leakage – in a chat setting, the model may have system or developer instructions and conversation history in its hidden context. Through multi-turn manipulation, users have managed to get models to divulge these hidden instructions or previous messages (a notable example was prompt leaks from ChatGPT revealing its formatting rules). These vulnerabilities are exacerbated over long interactions as more sensitive info accumulates in the context. Mitigating privacy leakage requires techniques like 1) reducing memorization during training (so the model generalizes rather than stores exact data points), 2) real-time filters to catch when a response may contain sensitive content, and 3) perhaps limiting the duration or content of contexts the model can access if they contain private data. As of now, completely preventing an LLM from regurgitating memorized secrets is an open problem – efforts are ongoing to quantify and bound such leakage [272] and to develop safer training regimes. In summary, multi-turn use of LLMs raises serious privacy considerations, and researchers must ensure that extended dialogues do not become a loophole for extracting confidential information or personal data.

### v5

@@ 4.2.2 Topic Switching & Discontinuous Reasoning
Real conversations frequently shift topics or revisit earlier threads after detours, posing significant challenges for LLMs. Models either fail to retain the prior thread when a topic is resumed or inappropriately carry over context from an unrelated topic—neither behavior matching human conversational fluency. Effective attention re-allocation across non-linear dialogue flows remains unsolved. Benchmarks such as MT-Bench (Zheng et al., 2023) and MT-Bench-101 (Bai et al., 2024) explicitly test category shifts, and even top-tier models frequently break coherence under non-linear conversation flows. Better mechanisms for tracking multiple concurrent topics and segmenting context are needed.
@@ 4.2.3 Proactive Information Seeking
In complex interactive settings such as medical diagnosis, troubleshooting, or tutoring, an ideal agent should proactively ask clarifying questions rather than passively answer. Most general-purpose LLMs, however, are largely reactive: trained primarily on single-turn Q&A, they lack exposure to multi-turn dialogue policies governing when to ask versus when to inform (Chen et al., 2025a), and tend to produce immediate answers even to underspecified queries. In a medical context, a user stating "I feel sick" should prompt targeted symptom questions; most LLMs instead offer generic advice without seeking clarification. Some recent work addresses this in constrained settings such as ambiguous text-to-SQL tasks (Chen et al., 2025a), but robust proactive dialogue behavior across arbitrary domains remains an open challenge.
@@ 4.2.4 Multilingual & Code-Switching Scenarios
Multi-turn interactions spanning multiple languages introduce additional complexity. While LLMs can handle single-turn non-English queries reasonably well, maintaining coherent cross-lingual context across turns is far more difficult. Code-switching—alternating languages across or within turns—often causes models to lose referential links or respond in the wrong language. Evaluation has been heavily skewed toward English, and performance degrades noticeably in lower-resource languages. There are also safety implications: mixing languages can circumvent content filters and expose weaknesses in multilingual alignment. Yoo et al. (2025) find that code-switching prompts can elicit undesirable behaviors that would not surface in English-only inputs, revealing gaps in multilingual robustness. Handling multi-turn dialogue in a culturally and linguistically consistent manner thus remains an open challenge.
@@ 4.3.1 Dynamic Preference & Objective Adaptation
Unlike human assistants, LLMs have no true long-term personalization within a conversation—they rely solely on the provided context window. While current chat models follow explicit style instructions, they often miss subtler cues or gradual shifts in user preferences. An ideal assistant might proactively simplify its language upon detecting user confusion; today’s LLMs typically require explicit instruction to do so. Dynamic adaptation is challenging because it requires inferring and retaining user preferences across turns without explicit reminders, yet model parameters are fixed during inference—any adaptation must emerge entirely from processing the conversation context. This leaves LLMs prone to either forgetting preferences or applying inferred styles inappropriately. Some initial work has explored updating pseudo-personas or user profiles during conversation, but reliable and safe mechanisms remain unresolved.
@@ 4.3.2 Knowledge Adaptation
Human conversationalists naturally absorb new facts and adjust to corrections during dialogue. Current LLMs, by contrast, have a fixed knowledge base at deployment: information provided mid-conversation resides only in the context window and is lost once the session resets. Enabling persistent memory and continual learning in interactive settings is an important open challenge (Wu et al., 2024a). Retrieval-augmented generation partially addresses this by fetching relevant facts from external databases, but still depends on pre-established knowledge sources. True on-the-fly learning—updating internal representations from user-provided data without retraining—remains difficult due to catastrophic forgetting and the risk that users could inject false or malicious information. Recent work has explored external memory modules and dynamic context augmentation as surrogates, writing new facts to a persistent scratchpad. Nonetheless, balancing plasticity and stability in real time remains unsolved.
@@ 4.3.3 Robustness to Misinformation & Adversarial Inputs
Multi-turn interactions introduce new attack surfaces: malicious users can gradually manipulate context or employ social-engineering tactics across turns to elicit unsafe outputs where single-turn attacks would fail. Recent studies confirm that even top-tier aligned models such as GPT-4 can be coerced into policy-violating behavior through carefully staged multi-turn strategies (Zhou and Arel, 2025). Beyond deliberate attacks, users may also introduce subtly incorrect facts across turns; lacking real-time fact-checking, LLMs tend to propagate such falsehoods. Prompt leakage—models inadvertently revealing hidden system instructions under sustained pressure—poses an additional concern. Collectively, these adversarial scenarios underscore that static safety training is insufficient when attacks are staged gradually. Building conversational agents capable of detecting inconsistent or malicious inputs and responding safely without disrupting the interaction remains an open research problem.
@@ 4.4.1 Scalable Data Curation
Achieving scalable data collection for multi-turn interactions remains a significant challenge at the intersection of data engineering and model training. Promising approaches include leveraging domain-specific real dialogues, forging institutional collaborations, and developing synthetic data pipelines—yet scalability in curating conversational datasets remains an open and pressing problem.
@@
As discussed in §2.1.1 and §2.2.2, healthcare uniquely benefits from access to extensive real-world patient-doctor conversational datasets that inherently provide rich diversity and realism, making them well-suited for continual pre-training, SFT, and RLHF. Replicating such collection pipelines in other domains is difficult: most fields lack established protocols for recording and anonymizing real-world conversations, and meeting rising demand for specialized conversational data will require broader collaboration among researchers, governments, and institutions.
Recent work has explored converting single-turn or limited-turn exchanges into richer multi-turn dialogues—either manually or via automated frameworks (Qiu et al., 2024; Li et al., 2023a; Wang et al., 2024d)—and using LLMs for synthetic generation and rewriting (Yang et al., 2024; Ding et al., 2023; Maheshwary et al., 2025; Xu et al., 2024). However, synthetic approaches carry well-documented drawbacks (Seddik et al., 2024; Chen et al., 2024e; c): they tend to lack the spontaneity and nuance of real conversations, risk amplifying factual inaccuracies and unnatural patterns, and over-rely on generic conversational templates. Going forward, progress requires not merely increasing data volume but improving quality—through better dialogue generation, robust filtering of synthetic outputs, and shared multi-domain repositories of multi-turn dialogues.
4.4.2 Metric Design
@@
Beyond datasets, the metrics for evaluating multi-turn interactions remain inadequate. Traditional measures such as single-turn accuracy or BLEU fail to capture dialogue-level qualities, motivating both fine-grained turn-level metrics and holistic conversation-level criteria.
Fine-Grained Capability Evaluation
@@
Benchmarks like MT-Bench-101 evaluate discrete abilities—logical consistency, factual recall, politeness—at the turn level (Bai et al., 2024), revealing which skills degrade over a conversation. Fine-grained scoring can expose, for instance, that a model maintains fluency but loses factual accuracy after many turns. Designing such rubrics is challenging, requiring careful identification and weighting of sub-skills along with expert annotation or LLM-as-judge evaluation. MultiChallenge’s instance-level rubrics (Deshpande et al., 2025) represent progress in this direction, but no universally adopted standard has yet emerged.
Long-Term Effectiveness
@@
Evaluating the global quality of a multi-turn exchange remains an open problem. Ideal metrics would capture whether the conversation as a whole succeeded—whether the user’s goal was achieved and the model remained helpful and coherent throughout. Some work measures conversational return-on-investment, assessing whether additional turns genuinely help or introduce confusion (Kwan et al., 2024), while others probe memory retention and consistency across turns. Metrics that directly quantify sustained performance over 10+ turns, or eventual convergence to a correct and useful outcome, remain underdeveloped. Evaluation measures analogous to task success rates in traditional dialogue systems are an important open research direction.
Cultural and Sociolinguistic Diversity
@@
Most evaluation setups remain Western and English-centric, limiting assessments of model performance for diverse users. A model may perform well on formal Q&A but poorly on casual, code-mixed, or culturally idiomatic conversations. FairMT-Bench (Fan et al., 2025) is a step toward addressing bias and fairness across demographic attributes in multi-turn settings, but defining metrics that quantify model behavior across a broad sociolinguistic spectrum—capturing bias, respectfulness, and user satisfaction for varied user profiles—remains an important and largely unsolved challenge.
@@
Multi-turn dialogue evaluation has increasingly turned to LLM-as-a-judge frameworks, where a strong model such as GPT-4 assesses response quality at scale. MT-Bench and Chatbot Arena exemplify this approach, achieving high consistency at low cost (Zheng et al., 2023). GPT-4-based evaluators align with human preferences roughly 80% of the time, approaching inter-annotator agreement levels (Zheng et al., 2023). However, AI judges exhibit systematic biases—including self-enhancement bias (favoring outputs from similar models) and verbosity bias (rewarding unnecessarily long responses)—and can miss subtle factual errors or nuanced value judgments that human reviewers would catch (Chen et al., 2024b). To mitigate these issues, researchers have proposed hybrid approaches combining LLM scoring with human oversight, such as prompting judges with explicit rubrics and incorporating periodic human audits, aiming to preserve the scalability of automated evaluation while maintaining reliability closer to human standards.
@@ 4.5.1 Bias Amplification
Multi-turn dialogues can inadvertently magnify model biases across turns: a biased assumption introduced early may be reinforced by subsequent exchanges, and the interactive nature of dialogue can cause the model to adapt to the user’s own biases, compounding problematic content. Fan et al. (2025) confirm this risk, finding that LLMs accumulate bias more readily in multi-turn settings than in single-turn prompts—a slight gender bias in an early response, for instance, can intensify as the conversation continues. FairMT-Bench reveals significant variability across state-of-the-art models, with many showing degraded fairness under back-and-forth discussions involving sensitive attributes. Addressing this challenge likely requires mechanisms beyond single-response detoxification, including dialogue-level self-monitoring, diverse data augmentation, and consistency enforcement with ethical guidelines across turns.
@@ 4.5.2 Privacy Leakage
Prolonged conversations increase the risk of LLMs revealing sensitive information—either disclosed by the user during dialogue or memorized from training data. Nasr et al. (2025) demonstrate that LLMs can memorize training data such as phone numbers and addresses, and that adversaries can extract this information through crafted prompt sequences. Multi-turn interactions amplify this threat by allowing iterative, context-building probes that may succeed where a single query would not. Beyond memorization, multi-turn manipulation has been used to elicit system prompt or context leakage, with users extracting hidden instructions or prior conversation contents. Mitigations include reducing memorization during training, deploying real-time filters for sensitive content, and restricting model access to private context. Completely preventing memorized-data leakage remains an open problem, with ongoing efforts to quantify its extent and develop safer training regimes (Nasr et al., 2025).